Find the vertex of the parabola formed by the following quadratic function: y = -x2 + 10x - 1
Answer by jrfrunner(365)


1. For ax^2 +bx+c parabola, the vertex (x,y) is such that x=-b/(2a)
for -x^2+10x-1, x=-10/(-2*(-1)) =5, y = -(5)^2+10*(5)-1=-25+50-1=24
vertex (5,24)
--
there are many other ways to get the vertex
* solve for the perfect square
* Graph
* take the derivative